OpenStudy (calculusxy):

Ramon has $4,200 to invest for college. b) Ramon's goal is to have $5,000 after 4 years. Is this possible if he invests with a return rate of 6%?

OpenStudy (calculusxy):

@perl

OpenStudy (calculusxy):

I have no idea of how to approach this problem.

OpenStudy (calculusxy):

@perl

OpenStudy (perl):

I think we can use the formula Balance = P( 1 + r*t)

OpenStudy (calculusxy):

Ok

OpenStudy (perl):

His goal is to get $5,000 after 4 years. Is this possible if he invests with a return rate of 6%? We can plug this into the equation B = P ( 1 + r * t ) where B = future balance P = principal r = rate of return t = # years This gives us: 5000 = P*( 1 + 6/100 * 4 )

OpenStudy (calculusxy):

I figured that p= 20833.33

OpenStudy (calculusxy):

So it is possible?

OpenStudy (perl):

oh wait there is missing information

OpenStudy (calculusxy):

So this is the whole problem: a. If Ramon invests $4,200 for 3 years and earns $630, what was the simple interest rate? b. (...)

OpenStudy (calculusxy):

Yes

OpenStudy (perl):

ok he invested 4200, that will be the principal

OpenStudy (perl):

4200*( 1 + 6/100 * 4 ) = amount at end of 4 years

OpenStudy (calculusxy):

1008

OpenStudy (perl):

that should equal to 5208

OpenStudy (calculusxy):

Okay. Thank you so much

OpenStudy (perl):

so now you can answer your question